OpenStudy (anonymous):

Simplify each radical expression. Sqrt 7(sqrt 14+sqrt 3)

OpenStudy (anonymous):

Can anyone pls help me on this problem thnx...

OpenStudy (espex):

Distribute your root7 in as you would any other similar expression.

OpenStudy (anonymous):

How u do that pls explain thnx..

OpenStudy (espex):

Do you know what the distributive property is?

OpenStudy (espex):

Essentially you "distribute" the multiplicand \(\sqrt{7}\) into the expression inside the parenthesis. So you will multiply both of the terms inside by the term outside.

OpenStudy (espex):

\(a(b+c)\rightarrow ab+ac\)

OpenStudy (anonymous):

I still don't get it..

OpenStudy (espex):

What exactly is it that is confusing you?

OpenStudy (espex):

It is not a minus sign, it is an arrow indicating a transition.

OpenStudy (anonymous):

Is my answer correct 17

OpenStudy (espex):

No, this will not give you an integer for an answer. Show me the process, step by step, how you got this number and I will show you where you went wrong.

OpenStudy (anonymous):

I'm using this calculator that's how I get the answer so if u could break it down for me ill do my best to understand it

OpenStudy (espex):

Using your calculator is why you are not understanding the problem. You have the expression \(\sqrt{7}(\sqrt{14}+\sqrt{3})\) what is your first step?

OpenStudy (anonymous):

7^2(14+3) is this right step

OpenStudy (espex):

Where did the \(7^2\) come from and what how did you get rid of the radical?

OpenStudy (anonymous):

Sorry I don't really understand the expression n radical thing if I could pls show me how pls thnx..

OpenStudy (espex):

Don't try and simplify it, just go through the steps first. What would you do first with \[\sqrt{7}(\sqrt{14}+\sqrt{3})\]

OpenStudy (anonymous):

Try to find the sqrt first?and add them together

OpenStudy (espex):

No, you distribute the \(\sqrt{7}\) in and multiply both the terms.

OpenStudy (anonymous):

^7(^14)+^7(3) like this?

OpenStudy (espex):

If by the carrot (^) you mean a root, yes. You multiply in to get \((\sqrt{7}*\sqrt{14})+(\sqrt{7}*\sqrt{3})\)

OpenStudy (espex):

Next you need to recall that \(\sqrt{a}*\sqrt{b}\) can be rewritten as \(\sqrt{a*b}\)

OpenStudy (anonymous):

^(7*14)=^(7*3). (98+21)=119

OpenStudy (espex):

You cannot combine the two as you did there because you are adding the roots of two numbers. What you will need to do is see if you can create squares that you can cancel and pull out from under the radical.

OpenStudy (anonymous):

Ok I'll try something else what abt =^98+3^7

OpenStudy (espex):

You will have \(\sqrt{7*14}+\sqrt{7*3}\) if you break it down into pieces to look for squares you need to turn them into their prime factors. If you notice, the second term here is already in prime factors. How would you break up the first term?

OpenStudy (anonymous):

^7*^14*^3 or ^98+3^7

OpenStudy (espex):

Prime factors are those that cannot be reduced further, what are the prime factors of 7 and 14?

OpenStudy (anonymous):

3 and 4?

OpenStudy (anonymous):

Prime factor of 7 is 1 and 7 prime factor of 14 is 4 and 14?

OpenStudy (anonymous):

R u still there?